# Efficient Reasoning

### v2

@@ Abstract
[…]
@@ 1 Introduction
Figure 1: Test-time scaling; the horizontal axis represents relative wall clock duration on an AMD MI300X GPU averaged across 553 GenEval Ghosh et al. [2023] prompts and the vertical axis represents the Overall GenEval score. VQAScore Lin et al. [2024] is used as verifier and FLUX-Schnell Labs et al. [2025] as generator for all methods. The Verifier-Threshold method is consistently more efficient and delivers higher scores compared to the baseline FlowRBF Kim et al. [2025] as well as other methods.
Image generation using AI has become a popular application with several models DeepMind [2025], OpenAI [2025], Apple Machine Learning Research [2025], Midjourney [2025], Stability AI [2024] being widely used. However, even state-of-the-art models often struggle to accurately follow detailed prompt instructions Ghosh et al. [2023], Huang et al. [2023]. To address this limitation, the vision community has drawn inspiration from the language domain by incorporating test-time compute Snell et al. [2024] and reasoning Guo et al. [2025], leading to the development of analogous approaches for vision tasks Ma et al. [2025], Zhuo et al. [2025]. However, reasoning-based methods inherently demand computational power and time during inference, which becomes costly at large scales Jin et al. [2025] or on resource-constrained devices Stelia [2025]. Consequently, improving the efficiency of these test-time compute algorithms becomes essential.
Fig. 1 illustrates how test-time scaling substantially increases computational requirements. The baseline implementation Kim et al. [2025] requires between $3\times$ and $10\times$ more compute time over regular image generation to achieve performance gains of $8\%$ to $12\%$ respectively on GenEval Ghosh et al. [2023]. In contrast, our proposed method, Verifier-Threshold, reaches the same $12\%$ improvement using just $(1/4)^{\text{th}}$ of the computational cost of the baseline, clearly demonstrating its efficiency. Furthermore, our algorithm continues to benefit from scaling, achieving gains of up to $15\%$.
The key contributions of this work are twofold: identifying the “compute-dumping” issue in the existing baseline algorithm and introducing the “Verifier-Threshold” algorithm as a solution. We validate our approach on state-of-the-art image generation models and datasets. We organize the rest of the paper as follows: Sec. 2 discusses the baseline algorithm and experimental setup. Sec. 3 shows the compute dumping phenomenon and possible solutions, including our proposed Verifier-Threshold approach. Lastly, Sec. 4 presents results, including some example images generated by the different approaches.
@@ 2 Preliminaries
2.1 Test-time Scaling and FlowRBF
Diffusion models Dhariwal and Nichol [2021], Ho et al. [2020] and flow models Labs et al. [2025] along with autoregressive models Tian et al. [2024] have become the mainstream methods for image generation. Test-time scaling for image generation basically aims to use the compute available at test-time to improve performance, either through increasing the denoising steps Salimans and Ho [2022] or searching for better noises Ma et al. [2025], Kim et al. [2025]. Noise-search in particular has shown to scale much better than denoising steps Ma et al. [2025]; hence we focus on that here. FlowRBF Kim et al. [2025] proposes a key idea of non-uniform inference-compute budget allocation across denoising steps which is a crucial advance in noise-search approaches. Using this idea, it beats all previous baseline methods such as Best-of-$N$ (which performs as good as zero-order and Search-over-Paths) from Ma et al. [2025]. Hence we use FlowRBF Kim et al. [2025] as our main baseline, and show improvements on the same flow models as proof-of-concept. However, FlowRBF’s approach of non-uniform allocation and our proposed algorithm can be easily extended to other diffusion models. Prior works are discussed in detail in Appendix B.
@@ 2.2 Experimental Setup
We use the FLUX-Schnell model Labs et al. [2025] for all our experiments similar to Kim et al. [2025]. We use ImageReward Xu et al. [2023] and VQAScore Lin et al. [2024] as our verifier reward models to guide the noise-search process. For evaluating the performance of different algorithms, we use the GenEval benchmark Ghosh et al. [2023], which consists of prompts that test for attributes such as color, number, position, etc. and helps determine an objective answer. We calculate the average percentage of correct images across different attributes as our performance metric. For the efficiency metric, we use the average wall clock duration (relative to base SDv2.1) taken for each prompt on a single AMD MI300X GPU. We also use the number of function evaluations (NFEs) to indicate the efficiency of the approach (lower is better). More details are discussed in Appendix A.
@@ 3.1 Compute Dumping and Manual Budgeting
The efficacy of the FlowRBF Kim et al. [2025] approach can be mainly attributed to the idea of non-uniform inference-compute allocation across denoising steps or rollover budget forcing (RBF). The basic concept behind the RBF algorithm is that the denoiser stays in a particular denoising step until either: (a) the verifier score for the new noise particle exceeds the previous step’s highest score, or (b) the per-step maximum compute budget is exhausted. The issue with these criteria is that: (a) it is greedy and jumps to the next step even with a marginally better particle, and (b) it does not account for the “wrong” trajectories the diffusion process takes due to the per-step budget.
Manifestations of these issues are shown in Figs. 2 and 3 where the model performs a ten-step denoising process. Fig. 2 (a, b, c) show that even though the RBF algorithm is designed to allow non-uniform allocation of inference-compute (NFEs), most of the steps (1 to 8) are allocated similar NFEs (except for the zeroth step which is hard-coded to use a pre-defined amount of NFEs). Significant portions of the total test-time NFEs are “dumped” at the last step. When the NFE budget is lower (=40) (Fig. 2 (a)), the algorithm quickly exhausts the per-step budget and moves to the next step. When the NFE budget is higher (=80, 160) (Figs. 2 (b & c)), the greediness of the RBF algorithm causes it to quickly move to the next step. In both cases, the algorithm has a large NFE budget remaining which it is forced to "dump" in the final denoising step. Fig. 2 (d) shows that the average verifier score at the last denoising step does not increase in proportion to the allocated compute, thus underscoring the dumping issue. Fig. 3 (a) shows an example prompt that highlights the mistake caused by the compute dumping phenomenon.
Intuitively, the RBF budget allocations in Fig. 2 are sub-optimal since the structural changes necessary to obtain a semantically correct image happens in the initial denoising steps, implying that more test-time budget needs to be allocated there rather than at the end. Manually hard-coding the compute distribution to have more budget at the initial steps is one approach to solving this problem as shown in Fig. 3 (b). Although this approach may yield improvements for certain prompts, it does not generalize well across all prompts. Therefore, an algorithm that automatically redistributes the compute budget is necessary.
@@
Figure 3: (a) shows three images generated at different denoising steps for the prompt “a photo of a bench left of a bear” using FlowRBF Kim et al. [2025], with a total of 40 NFEs and ImageReward Xu et al. [2023] as the verifier. This example illustrates the compute-dumping issue. The first image (at the second denoising step) captures the correct structure, though the texture remains noisy. However, the second image, produced at the third step, achieves the highest verifier score and is therefore selected, despite placing the bear on top of the bench rather than to the left. Because no additional budget is allocated at earlier steps, the denoising process continues from this incorrect latent, ultimately producing a final image with the bear on top of the bench. (b) demonstrates how this issue could be mitigated by manually reallocating the compute budget through hard-coded NFEs at each step (i.e., manual budgeting).
@@ 3.2 Verifier-Threshold Algorithm
Rather than advancing to the next denoising step as soon as a higher-reward noise particle is obtained, we propose waiting until the verifier score surpasses the previous one by a significant margin (i.e., a threshold). In this way, the algorithm progresses only when the score at a given step is sufficiently improved. This mechanism forms the basis of our Verifier-Threshold algorithm, as illustrated in Fig. 4a and detailed in Alg. 1.
@@ 4 Results
As described in Sec. 2, we evaluate the performance of different algorithms on the GenEval benchmark Ghosh et al. [2023]. Fig. 1 highlights the advantages of our approach when using VQAScore Lin et al. [2024] as the verifier. To demonstrate that these benefits generalize across verifiers, we also evaluate performance using ImageReward Xu et al. [2023], as shown in Fig. 5. In both cases, our method achieves $2\times$ to $4\times$ higher efficiency for the same benchmark score, or alternatively, a $4\%$ to $5\%$ improvement in benchmark score at the same computational cost. All reported results correspond to test-time compute within a practical range of 2–3 minutes. Additionally, Fig. 6 presents example prompts where our algorithm produces objectively correct outputs compared to baseline methods. More comprehensive results and further examples are provided in Appendix C.
Figure 5: Test-time scaling when ImageReward Xu et al. [2023] is used as verifier.
Figure 6: The outputs of the model for the prompt “a photo of a wine glass right of a hot dog”. (a), (b), (c) represent the images generated by FLUX-Schnell Labs et al. [2025] with no test-time scaling, FlowRBF Kim et al. [2025], and Verifier-Threshold, respectively. Figures (a) & (b) have the wine glass on the left while (c) accurately provides what the prompt is requesting.
@@ 5 Conclusion
In this work, we identify a key flaw (compute dumping) in the state-of-the-art test-time scaling algorithm Kim et al. [2025] and propose the Verifier-Threshold algorithm as an effective solution. We demonstrate its performance and efficiency advantages across multiple verifiers on the GenEval benchmark. For future work, we plan to investigate adaptive verifier thresholds that vary across denoising steps and explore strategies to automate their selection. We also aim to extend our approach to standard diffusion models (beyond flow-based models).
@@ A.1 Dataset & Models
For evaluating our models we employ the widely-used GenEval dataset Ghosh et al. [2023]. It contains 553 prompts which checks six different attributes: single object, two object, counting, colors, position, and attribute binding. For the generator, we use the FLUX-Schnell model Labs et al. [2025] extensively due to its SOTA performance and baseline implementation from FlowRBF Kim et al. [2025]. We run it consistently for ten denoising steps for best results. For the verifiers, we use ImageReward Xu et al. [2023] and VQAScore Lin et al. [2024] which are both variants of the CLIP model Radford et al. [2021].
A.2 Methods
We use the same setup and hyperparameters as FlowRBF Kim et al. [2025] for all the methods: Best-of-$N$, search-over-paths, SVDD, and FlowRBF itself. We use a verifier threshold value ($\Delta$) of 0.005 and 0.00125 for ImageReward and VQAScore respectively when the total NFEs is 40. For subsequent total NFEs of 80 and 160, we correspondingly multiplied the threshold values with $2\times,4\times$ the above value.
Appendix B Related Work
@@
In recent years, scaling of AI models has mainly involved increasing model size and the amount of compute and data utilized during training. However, recent work shows that spending more compute at test-time can substantially improve output quality and reasoning capability, thus creating a new dimension for scaling compute to improve model performance. OpenAI’s o1 series popularized this paradigm by spending more test-time budget on deliberate “thinking” before answering OpenAI [2024]. DeepSeek-R1/R1-Zero similarly leverages extended reasoning trajectories at inference Guo et al. [2025]. Snell et al. formalize when scaling test-time compute can be more effective than scaling model size Snell et al. [2024]. Complementing these analyses, s1 demonstrates a simple and effective test-time scaling recipe with strong empirical gains Muennighoff et al. [2025].
@@
Reward-based sampling methods repeatedly sample from the model’s learned distribution, guided by a reward such as text–image alignment or aesthetic quality, to select higher-quality outputs. The most basic example is Best-of-$N$ (BoN), which generates $N$ independent images with different random seeds and picks the one with the highest reward. Ma et al. (2025) formalize this paradigm by introducing both BoN and Search-over-Paths (SoP), the latter refining trajectories by sampling particles along the denoising path Ma et al. [2025]. Building on this idea, “noise trajectory search” methods incorporate reward-guided selection directly into the denoising process: SVDD Li et al. [2024a] selects, at every denoising step, the particle with the highest expected reward. Extending reward-based search to flow models, Kim et al. (2025) Kim et al. [2025] propose three complementary techniques: test-time SDE conversion, which introduces stochasticity into otherwise deterministic flows; interpolant conversion, which alters the trajectory interpolant to expand the search space; and Rollover Budget Forcing (RBF), which dynamically reallocates compute across timesteps by advancing any particle that exceeds the previous best reward and rolling over unused function evaluations to future steps, thereby ensuring efficient utilization of the total budget.
Appendix C Comprehensive Results
In this section, we flush out the plots in Fig. 1 & Fig. 5 in the form of Tab. 1. We also show the images generated for various example prompts, especially in cases where our approach succeeds when other baselines fail, and also some cases where all approaches fail. In all the example images, Fig. (a) corresponds to the image generated without test-time scaling, Fig. (b) corresponds to the one generated using FlowRBF Kim et al. [2025] and Fig. (c) is the one generated by our Verifier-Threshold algorithm.
@@ Appendix C Comprehensive Results
Steering Singhal et al. [2025] SDv2.1 ImageReward Xu et al. [2023] 4 3.57 57.93
Steering Singhal et al. [2025] SDv2.1 ImageReward Xu et al. [2023] 20 15.62 64.91
Steering Singhal et al. [2025] SDv2.1 ImageReward Xu et al. [2023] 40 31.91 66.18
BoN Ma et al. [2025] FLUX-Schnell ImageReward Xu et al. [2023] 40 9.70 70.62
BoN Ma et al. [2025] FLUX-Schnell ImageReward Xu et al. [2023] 80 17.16 70.36
BoN Ma et al. [2025] FLUX-Schnell ImageReward Xu et al. [2023] 160 26.39 72.45
BoN Ma et al. [2025] FLUX-Schnell VQAScore Lin et al. [2024] 40 10.29 70.07
BoN Ma et al. [2025] FLUX-Schnell VQAScore Lin et al. [2024] 80 18.72 70.54
BoN Ma et al. [2025] FLUX-Schnell VQAScore Lin et al. [2024] 160 27.07 71.91
SoP Ma et al. [2025] FLUX-Schnell ImageReward Xu et al. [2023] 250 21.00 60.68
SoP Ma et al. [2025] FLUX-Schnell VQAScore Lin et al. [2024] 250 22.25 61.19
SVDD Li et al. [2024b] FLUX-Schnell ImageReward Xu et al. [2023] 250 43.69 68.76
SVDD Li et al. [2024b] FLUX-Schnell VQAScore Lin et al. [2024] 250 56.74 72.89
FlowRBF Kim et al. [2025] FLUX-Schnell ImageReward Xu et al. [2023] 40 12.80 70.10
FlowRBF Kim et al. [2025] FLUX-Schnell ImageReward Xu et al. [2023] 80 18.11 72.45
FlowRBF Kim et al. [2025] FLUX-Schnell ImageReward Xu et al. [2023] 160 34.11 74.20
FlowRBF Kim et al. [2025] FLUX-Schnell VQAScore Lin et al. [2024] 40 13.83 70.32
FlowRBF Kim et al. [2025] FLUX-Schnell VQAScore Lin et al. [2024] 80 27.03 71.77
FlowRBF Kim et al. [2025] FLUX-Schnell VQAScore Lin et al. [2024] 160 48.98 73.45
VT (Ours) FLUX-Schnell ImageReward Xu et al. [2023] 40 10.31 71.61
VT (Ours) FLUX-Schnell ImageReward Xu et al. [2023] 80 18.68 76.77
VT (Ours) FLUX-Schnell ImageReward Xu et al. [2023] 160 39.36 78.62
VT (Ours) FLUX-Schnell VQAScore Lin et al. [2024] 40 12.80 73.83
VT (Ours) FLUX-Schnell VQAScore Lin et al. [2024] 80 25.90 76.12
VT (Ours) FLUX-Schnell VQAScore Lin et al. [2024] 160 51.97 77.20
Table 1: Comparison of different approaches across models, verifiers, NFEs, along with their efficiency (wall-clock time) & performance (GenEval) scores. Methods are Best-of-NN (BoN), Search-over-Paths (SoP), Soft Value-based Decoding in Diffusion (SVDD), and Verifier-Threshold (VT).
Figure 7: “a photo of a purple wine glass and a black apple”. For all examples in this section, Fig. (a) is the image generated without test-time scaling, Fig. (b) is generated using FlowRBF Kim et al. [2025] and Fig. (c) is generated by our Verifier-Threshold algorithm.

### v3

# Verifier Threshold: An Efficient Test-Time Scaling Approach for Image Generation

@@ Abstract
Image generation has emerged as a mainstream application of large generative models. Just as test-time compute and reasoning have improved language model capabilities, similar benefits have been observed for image generation models. In particular, searching over noise samples for diffusion and flow models has been shown to scale well with test-time compute. While recent works explore allocating non-uniform inference-compute budgets across denoising steps, existing approaches rely on greedy heuristics and often allocate the compute budget ineffectively. In this work, we study this problem and propose a simple fix. We propose Verifier-Threshold, which automatically reallocates test-time compute and delivers substantial efficiency improvements. For the same performance on the GenEval benchmark, we achieve a $2\times$- $4\times$ reduction in computational time over the state-of-the-art method.
@@ 1 Introduction
Figure 1: Test-time scaling; the horizontal axis represents relative wall-clock duration on an AMD MI300X GPU averaged across 553 GenEval (Ghosh et al., 2023) prompts and the vertical axis represents the overall GenEval score. VQAScore (Lin et al., 2024) is used as the verifier and FLUX-Schnell (Labs et al., 2025) as the generator for all methods. Verifier-Threshold is consistently more efficient and delivers higher scores than the baseline FlowRBF and other methods.
Image generation has become a popular application of AI, with several models (DeepMind, 2025; OpenAI, 2025; Apple Machine Learning Research, 2025; Midjourney, 2025; Stability AI, 2024) now widely used. However, even state-of-the-art models often struggle to accurately follow detailed prompt instructions (Ghosh et al., 2023; Huang et al., 2023). To address this limitation, the vision community has drawn inspiration from language models by incorporating test-time compute (Snell et al., 2024) and reasoning (Guo et al., 2025), leading to the development of analogous approaches for vision tasks (Ma et al., 2025; Zhuo et al., 2025). Reasoning-based methods, however, inherently demand substantial computation and wall-clock time at inference, which becomes costly at scale (Jin et al., 2025) or on resource-constrained devices (Stelia, 2025). Consequently, improving the efficiency of test-time compute algorithms is essential.
Fig. 1 illustrates how test-time scaling increases computational requirements. The baseline implementation (Kim et al., 2025) requires between $3\times$ and $10\times$ more compute time than regular image generation to achieve GenEval (Ghosh et al., 2023) gains of $8\%$ to $12\%$, respectively. In contrast, our proposed method, Verifier-Threshold, reaches the same $12\%$ improvement using only $(1/4)^{\text{th}}$ of the baseline’s computational cost, clearly demonstrating its efficiency. Furthermore, our algorithm continues to benefit from scaling, achieving gains of up to $15\%$.
The key contributions of this work are twofold: identifying the “compute-dumping” issue in an existing baseline algorithm and introducing the Verifier-Threshold algorithm as a solution. We validate our approach on state-of-the-art image generation models and datasets. We organize the rest of the paper as follows: Sec. 2 discusses the baseline algorithm and experimental setup. Sec. 3 analyzes the compute dumping phenomenon and presents solutions, including Verifier-Threshold. Lastly, Sec. 4 presents results, including example images generated by the different approaches.
@@ 2 Preliminaries
2.1 Related Works
Diffusion models (Dhariwal and Nichol, 2021; Ho et al., 2020) and flow models (Labs et al., 2025) along with autoregressive models (Tian et al., 2024) have become the mainstream methods for image generation. Test-time scaling for image generation basically aims to use the compute available at test-time to improve performance, either through increasing the denoising steps (Salimans and Ho, 2022) or searching for better noises (Ma et al., 2025; Kim et al., 2025). Noise-search in particular has been shown to scale much better than increasing denoising steps (Ma et al., 2025); hence we focus on it here. FlowRBF introduces the key idea of non-uniform inference-compute budget allocation across denoising steps, which is a crucial advance in noise-search approaches. Using this idea, it outperforms prior baselines such as Best-of-$N$ (which performs as well as zero-order and Search-over-Paths) from (Ma et al., 2025). Hence, we use FlowRBF as our main baseline and show improvements on the same flow models as a proof of concept. FlowRBF’s approach of non-uniform allocation and our proposed algorithm can also be extended to other diffusion models. Prior work is discussed in detail in Appendix B.
@@ 2.2 Experimental Setup
We use the FLUX-Schnell model (Labs et al., 2025) for all experiments, following (Kim et al., 2025). We use ImageReward (Xu et al., 2023) and VQAScore (Lin et al., 2024) as verifier reward models to guide the noise-search process. Only open-source models are used throughout since these experiments cannot be conducted with closed-source models. To evaluate performance, we use the GenEval benchmark (Ghosh et al., 2023), which consists of prompts that test for attributes such as color, number, and position and provides objective answers. We report the average percentage of correct images across attributes as our performance metric. For efficiency, we use the average wall-clock duration per prompt on a single AMD MI300X GPU (reported relative to SDv2.1 which takes the least amount of time; see Appendix C). We also report the number of function evaluations (NFEs), where lower is better. More details are provided in Appendix A.
@@ 3.1 Compute Dumping and Manual Budgeting
The efficacy of FlowRBF can be mainly attributed to its non-uniform inference-compute allocation across denoising steps, termed rollover budget forcing (RBF). The basic concept behind RBF is that the denoiser remains at a denoising step until either: (a) the verifier score for a new noise particle exceeds the previous step’s best score, or (b) the per-step maximum compute budget is exhausted. The issue with these criteria is that (a) the method is greedy and advances even for marginal improvements, and (b) it does not account for “wrong” trajectories induced by per-step budget constraints.
Manifestations of these issues are shown in Figs. 2 and 3, where the model runs a ten-step denoising process. Fig. 2(a–c) show that although RBF is designed to enable non-uniform allocation of inference compute (NFEs), most steps (1–8) receive similar NFEs (except for step 0, which is hard-coded to use a predefined NFE). As a result, a large fraction of the test-time NFEs are “dumped” into the final step. When the NFE budget is lower (=40=40; Fig. 2(a)), the algorithm quickly exhausts the per-step budget and advances to the next step. When the budget is higher (=80,160=80,160; Fig. 2(b,c)), the greediness of RBF causes it to advance quickly to the next step as well. In both cases, substantial budget remains and is forced into the last denoising step. Fig. 2(d) shows that the average verifier score at the last denoising step does not increase proportionally with the allocated compute, underscoring the dumping issue. Fig. 3(a) shows an example prompt that highlights the resulting failure mode.
Intuitively, the RBF budget allocations in Fig. 2 are sub-optimal because the structural changes needed to obtain a semantically correct image often occur in early denoising steps, suggesting that more budget should be allocated early rather than at the end. Manually hard-coding the compute distribution to allocate more budget to initial steps is one way to address this issue (Fig. 3(b)). Although this can improve some prompts, it does not generalize well. Therefore, an algorithm that automatically redistributes compute is needed.
@@
Figure 3: (a) shows three images generated at different denoising steps for the prompt “a photo of a bench left of a bear” using FlowRBF, with a total of 40 NFEs and ImageReward (Xu et al., 2023) as the verifier. This example illustrates the compute-dumping issue. The first image (at the second denoising step) captures the correct structure, though the texture remains noisy. However, the second image, produced at the third step, achieves the highest verifier score and is therefore selected, despite placing the bear on top of the bench rather than to the left. Because no additional budget is allocated at earlier steps, the denoising process continues from this incorrect latent, ultimately producing a final image with the bear on top of the bench. (b) demonstrates how this issue could be mitigated by manually reallocating the compute budget through hard-coded NFEs at each step (i.e., manual budgeting).
@@ 3.2 Verifier-Threshold Algorithm
Rather than advancing to the next denoising step as soon as a higher-reward noise particle is obtained, we propose waiting until the verifier score surpasses the previous one by a significant margin (i.e., a threshold). In this way, the algorithm progresses only when the score at a given step improves sufficiently. This mechanism forms the basis of our Verifier-Threshold algorithm, as illustrated in Fig. 4(a) and detailed in Alg. 1.
@@ 4 Results
As described in Sec. 2, we evaluate different algorithms on the GenEval benchmark (Ghosh et al., 2023). Fig. 1 highlights the advantages of our approach when using VQAScore (Lin et al., 2024) as the verifier. The choice of threshold value $\Delta$ is described in Appendix A. To show that these benefits generalize across verifiers, we also evaluate ImageReward (Xu et al., 2023), as shown in Fig. 5. In both cases, our method achieves $2\times$ to $4\times$ higher efficiency for the same benchmark score, or alternatively a $4\%$ to $5\%$ improvement in benchmark score at the same computational cost. All reported results correspond to practical test-time compute budgets (2–3 minutes). Additionally, Fig. 6 presents example prompts where our algorithm produces objectively correct outputs compared to baseline methods. More comprehensive results and further examples are provided in Appendix A.
Figure 5: Test-time scaling when ImageReward (Xu et al., 2023) is used as verifier.
Figure 6: The outputs of the model for the prompt “a photo of a wine glass right of a hot dog”. (a), (b), (c) represent the images generated by FLUX-Schnell (Labs et al., 2025) with no test-time scaling, FlowRBF, and Verifier-Threshold, respectively. Figures (a) & (b) have the wine glass on the left while (c) accurately provides what the prompt is requesting.
@@ 5 Conclusion
In this work, we identify a key flaw (compute dumping) in the state-of-the-art test-time scaling algorithm (Kim et al., 2025) and propose the Verifier-Threshold algorithm as an effective solution. We demonstrate its performance and efficiency advantages across multiple verifiers on the GenEval benchmark. For future work, we plan to investigate adaptive verifier thresholds that vary across denoising steps and explore strategies to automate their selection. There is also a lack of theoretical or interpretability grounding for test-time compute in the image generation community currently, and we will look into that as well as a future direction for our work.
@@ A.1 Dataset & Models
For evaluating our models we employ the widely-used GenEval dataset (Ghosh et al., 2023). It contains 553 prompts which checks six different attributes: single object, two object, counting, colors, position, and attribute binding. For the generator, we use the FLUX-Schnell model (Labs et al., 2025) extensively due to its SOTA performance and baseline implementation from FlowRBF. We run it consistently for ten denoising steps for best results. For the verifiers, we use ImageReward (Xu et al., 2023) and VQAScore (Lin et al., 2024) which are both variants of the CLIP model (Radford et al., 2021).
A.2 Hyperparameters & Verifier-Threshold Choice
We use the same setup and hyperparameters as FlowRBF for all the methods: Best-of-$N$, search-over-paths, SVDD, FlowRBF and ours. When using 40 NFEs, we used 5 initial noise samples, and doubled/quadrupled it for 80/160 respectively. We used guidance-scale 3.5, diffusion norm 3.0, exponential time-scheduler with $t_{\text{max}}=1000$ and image-size 1024 $\times$ 1024.
The threshold value $\Delta$ is chosen independently for each verifier. The difference between the average verifier score for the initial timestep and final timestep is divided by the number of timesteps to obtain it. We use a verifier threshold value ($\Delta$) of 0.005 and 0.00125 for ImageReward and VQAScore respectively when the total NFEs is 40. For subsequent total NFEs of 80 and 160, we correspondingly multiplied the threshold values with $2\times,4\times$ the above value. Fig. 7 shows the verifier score variation for ImageReward which was obtained when running a simple evaluation of the FlowRBF algorithm (NFE=80=80) for GenEval prompts. From the explanation above, we can calculate the value as $\Delta=\frac{1.62-1.52}{10}=0.01=2\times 0.005$.
Figure 7: Verifier score variation for ImageReward used to derive $\Delta$.
Appendix B Related Works
@@
In recent years, scaling of AI models has mainly involved increasing model size and the amount of compute and data utilized during training. However, recent work shows that spending more compute at test-time can substantially improve output quality and reasoning capability, thus creating a new dimension for scaling compute to improve model performance. OpenAI’s o1 series popularized this paradigm by spending more test-time budget on deliberate “thinking” before answering (OpenAI, 2024). DeepSeek-R1/R1-Zero similarly leverages extended reasoning trajectories at inference (Guo et al., 2025). Snell et al. formalize when scaling test-time compute can be more effective than scaling model size (Snell et al., 2024). Complementing these analyses, s1 demonstrates a simple and effective test-time scaling recipe with strong empirical gains (Muennighoff et al., 2025).
@@
Reward-based sampling methods repeatedly sample from the model’s learned distribution, guided by a reward such as text–image alignment or aesthetic quality, to select higher-quality outputs. The most basic example is Best-of-$N$ (BoN), which generates $N$ independent images with different random seeds and picks the one with the highest reward. Ma et al. (2025) formalize this paradigm by introducing both BoN and Search-over-Paths (SoP), the latter refining trajectories by sampling particles along the denoising path (Ma et al., 2025). Building on this idea, “noise trajectory search” methods incorporate reward-guided selection directly into the denoising process: SVDD (Li et al., 2024b) selects, at every denoising step, the particle with the highest expected reward. Extending reward-based search to flow models, Kim et al. (2025) (Kim et al., 2025) propose three complementary techniques: test-time SDE conversion, which introduces stochasticity into otherwise deterministic flows; interpolant conversion, which alters the trajectory interpolant to expand the search space; and Rollover Budget Forcing (RBF), which dynamically reallocates compute across timesteps by advancing any particle that exceeds the previous best reward and rolling over unused function evaluations to future steps, thereby ensuring efficient utilization of the total budget.
@@
Steering SDv2.1 ImageReward 4 3.57 57.93
Steering SDv2.1 ImageReward 20 15.62 64.91
Steering SDv2.1 ImageReward 40 31.91 66.18
BoN FLUX-Schnell ImageReward 40 9.70 70.62
BoN FLUX-Schnell ImageReward 80 17.16 70.36
BoN FLUX-Schnell ImageReward 160 26.39 72.45
BoN FLUX-Schnell VQAScore 40 10.29 70.07
BoN FLUX-Schnell VQAScore 80 18.72 70.54
BoN FLUX-Schnell VQAScore 160 27.07 71.91
SoP FLUX-Schnell ImageReward 250 21.00 60.68
SoP FLUX-Schnell VQAScore 250 22.25 61.19
SVDD FLUX-Schnell ImageReward 250 43.69 68.76
SVDD FLUX-Schnell VQAScore 250 56.74 72.89
FlowRBF FLUX-Schnell ImageReward 40 12.80 70.10
FlowRBF FLUX-Schnell ImageReward 80 18.11 72.45
FlowRBF FLUX-Schnell ImageReward 160 34.11 74.20
FlowRBF FLUX-Schnell VQAScore 40 13.83 70.32
FlowRBF FLUX-Schnell VQAScore 80 27.03 71.77
FlowRBF FLUX-Schnell VQAScore 160 48.98 73.45
VT (Ours) FLUX-Schnell ImageReward 40 10.31 71.61
VT (Ours) FLUX-Schnell ImageReward 80 18.68 76.77
VT (Ours) FLUX-Schnell ImageReward 160 39.36 78.62
VT (Ours) FLUX-Schnell VQAScore 40 12.80 73.83
VT (Ours) FLUX-Schnell VQAScore 80 25.90 76.12
VT (Ours) FLUX-Schnell VQAScore 160 51.97 77.20
Table 1: Comparison of different approaches across models, verifiers, NFEs, along with their efficiency (wall-clock time) & performance (GenEval) scores. Methods are Steering (Singhal et al., 2025), Best-of-NN (Ma et al., 2025) (BoN), Search-over-Paths (Ma et al., 2025) (SoP), Soft Value-based Decoding in Diffusion (Li et al., 2024a) (SVDD), FlowRBF (Kim et al., 2025), and Verifier-Threshold (VT). Verifiers are ImageReward (Xu et al., 2023) and VQAScore (Lin et al., 2024).
Appendix C Comprehensive Results
In this section, we expand the plots in Fig. 1 and Fig. 5 into Tab. 1. We also show images for various example prompts, including cases where our approach succeeds while other baselines fail, as well as cases where all approaches fail. In all example triplets, Fig. (a) corresponds to generation without test-time scaling, Fig. (b) corresponds to FlowRBF, and Fig. (c) corresponds to our Verifier-Threshold algorithm.
Figure 8: “a photo of a purple wine glass and a black apple”. For all examples in this section, Fig. (a) is the image generated without test-time scaling, Fig. (b) is generated using FlowRBF and Fig. (c) is generated by our Verifier-Threshold algorithm.